# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v2

@@ 5.2 Attack Success Rate in Causing Hazards
We consider an attack successful if a collision event is observed (Ego vehicle collides with the lead vehicle) or the relative distance between the lead vehicle and the Ego vehicle is no larger than 0m. The attack success rate across all 1000 simulations is reported, if not specified otherwise.
Figure 9: Top: Adversarial patch examples generated using the optimization-based method vs. random-based and FGSM-based methods. Bottom: Success rate of CA-Opt and baseline attacks in absence of safety interventions.
Fig. 9 shows the success rates for each attack. The CA-Random attack causes hazards in less than 1.2% of each testing scenario, resulting in a total success rate of 0.7%. This indicates that a randomly generated adversarial patch does not significantly alter the DNN model predictions, thereby causing few hazards in most simulations. We also observe that increasing the random perturbation values in the adversarial patch does not noticeably increase the success rate. This is because the DNN model in OpenPilot is trained to detect the presence of any front objects rather than predict their specific class. As a result, it exhibits better resilience to adversarial attacks. In addition, the generated patches with large perturbation values are much more visible to the driver compared to those generated by the optimization-based method (see Fig. 9 (Top)). Such attacks would attract the driver’s attention and might be prevented before causing any hazards. On the other hand, the CA-FGSM attack achieves an average success rate of 16.2%, but the generated patches are still noticeable and can be mitigated by the driver (see more details in Section 5.3 and Appendix D).
In contrast, the proposed CA-Opt attack achieves a 100% success rate in all testing scenarios, surpassing CA-Random by 142.9 times. The proposed algorithm for real-time updating of attack vectors (see Section 3.3.2) also enhances the attack success rates. We also observe that the Random-Opt baseline only achieves a success rate of 3.5% on average, 28.6 times lower than the CA-Opt, as it wastes resources by injecting perturbations at non-hazardous system states. This indicates the insufficiency of optimization-based methods alone in successfully attacking ACC at runtime and highlights the importance of strategic timing of attacks.
Observation 1: CA-Opt is more efficient than baselines in identifying the most critical times and optimal DNN perturbation values for attacking the ACC systems.
5.3 Attack Stealthiness with Safety Interventions
@@
5.3.1 Stealthiness in Perception Input
@@
Table V presents the results averaged over all the test scenarios and simulations. The CA-Opt attack achieves at least a 99.2% success rate under all three stealthiness levels and keeps the perturbation degree less than 0.015 ($L_{\infty}$) and 0.184 ($L_{2}$). The perturbed image with the adversarial patch has a similarity of UIQ = 0.993 (1 means identical) to the original image. We choose the $\lambda$ value to be $10^{-3}$ in our evaluations because of its stealthiness and high attack effectiveness. Examples of the generated adversarial patches (with $\lambda=10^{-3}$) are presented in Fig. 5 (see the zoomed-in area) and Fig. 9, which are almost invisible to human eyes.
To further evaluate the stealthiness of our attack design, we also conduct a user study with 30 participants. Results show that adversarial patches at $\lambda=10^{-2}$ and $\lambda=10^{-3}$ are almost imperceptible to human drivers, and the patches generated by CA-Opt attacks are less noticeable than those generated by baseline perception attacks (CA-Random and CA-FGSM) (refer to Appendix D for more details).
TABLE V: Attack success rate with different patch stealthiness levels $\lambda$ and pixel perturbation degrees.
@@
5.3.2 Evading Safety Interventions
For a more realistic evaluation of the effectiveness of different attack strategies, we rerun our experiments with different safety interventions (introduced in Section 4). A calibration of the safety features is performed before our evaluations to make sure the interventions are triggered correctly without any false positives.
We test each attack method with different AEBS configurations: (i) FCW/AEB depends on an independent camera that is not compromised, (ii) AEB/FCW utilizes compromised camera inputs similar to the ACC (simulating stock ACC and AEBS that share a camera or independent ACC and AEBS cameras that are both compromised), or (iii) AEB/FCW is disabled. Driver intervention and ACC safety constraint checking (OpenPilot Panda checks) are considered for all three settings. Here, we do not test the Random-Opt attack due to its similarity to CA-Opt but with worse performance. We assess the efficacy of each attack method using metrics such as the attack success rate, safety intervention activation rate (indicating the percentage of simulations triggering safety interventions), and hazard prevention rate (the percentage of simulations where hazards would occur without safety interventions).
Table VI shows the experimental results of each attack method with different safety intervention configurations. We observe that, regardless of the intervention configuration, the CA-Random and CA-FGSM attacks fail to cause any hazards due to their low baseline success rates (see Fig. 9) and their noticeable perturbations that trigger the driver interventions in 23.8-27.4% and 100% of scenarios. These findings highlight the effectiveness of human drivers in preventing accidents and keeping autonomous driving safe.
In contrast, with AEBS disabled (the last row of Table VI), the proposed CA-Opt attack successfully evades driver intervention in all simulations. This is attributed to the small, virtually invisible perturbation value of the adversarial patch, resulting in an average attack success rate of 100%. We also conduct experiments that simulate higher driver sensitivity levels by decreasing the mean perturbation value threshold for activating driver intervention from the default value of 15% (see Section 6) to 10%, 5%, 2%, 1.5%, 1%, and 0.5%. As shown in Fig. 10, when perturbation thresholds are set to 0.5% and 1% (representing highly sensitive driver), the CA-Opt attack triggers driver interventions in all and 79.6% of the simulations, leading to attack success rates of 0% and 20.4%, respectively. However, with thresholds higher than 1.5%, the proposed attack maintains a 100% success rate without triggering driver intervention. This finding underscores the robustness of the attack in evading driver detection across a range of driver sensitivities.
TABLE VI: Performance of attacks with all the safety features and different AEBS settings.
@@ 5.3.2 Evading Safety Interventions
5.3.3 Stealthiness in Control Output
The stealthy perturbations on the perception input can get propagated through the DNN model and ACC logic and lead to changes in ACC control output. Although the attacker’s goal is to cause sudden accelerations on the ACC output, large deviations in vehicle states may be detected by the human driver or existing safety and defense mechanisms.
To further evaluate the stealthiness of our proposed perception attack (CA-Opt), we examine deviations in autonomous vehicle states and control outputs resulting from the attack compared to two baseline output attacks, called MaxOut and StrategicOut. These attacks directly modify ACC output control commands, by setting them to a maximum allowed acceleration value (MaxOut) or a strategic value (StrategicOut) based on a method from prior research [17]. But they use the same context-aware method as CA-Opt for selecting the attack times and durations.
@@ 5.3.3 Stealthiness in Control Output
TABLE VII: Performance of StrategicOut attack with all the safety features and different AEBS settings.
@@
We also evaluate the success rate of StrategicOut’s attack under different safety intervention configurations presented in Section 5.3.2. We do not assess the MaxOut attack due to its high likelihood of being detected. Results in Table VII show that without AEBS, StrategicOut achieves a high success rate of 99.5% by generating attack values within safety limits and avoiding driver intervention. However, with AEBS active, using the same camera inputs as ACC, the success rate drops significantly to 20.3%, primarily due to safety interventions triggered in all simulations. No impact is observed when AEBS uses different camera inputs, as the StrategicOut attack does not affect the camera images.
Observation 4: CA-Opt perception attacks have advantages over direct output attacks in minimizing vehicle state changes to evade detection and safety interventions while maintaining high effectiveness in causing hazards.
6 Evaluation in Real-World Settings
The evaluations in this section answer the following research questions about the effectiveness of our attack in real-world settings.
RQ3: Is our attack robust to real-world factors such as different camera positions and weather conditions?
RQ4: Can our attack transfer well from simulation to real-world implementation?
RQ5: How does our attack affect the real-time system operation?
RQ6: Can our attack evade detection or mitigation by the existing adversarial patch defense methods?
6.1 Robustness to Real-world Factors
@@
Fig. 12 illustrates the 100% success rate of our CA-Opt attack across 12 testing scenarios. The Ego vehicle initially maintains a safe following distance, deviates from it around the 2,500-3,000 control cycle or step due to the adversarial patch, and collides with the lead vehicle. These experimental results demonstrate our attack is robust to different camera positions and initial longitudinal distances and can cause severe adverse events (e.g., collision).
@@
Figure 12: Relative distance trajectories under CA-Opt attack with different camera heights (H1:1.1m, H2:1.3m, H3:1.5m, H4:1.7m) and initial longitudinal distances between the lead vehicle and the Ego vehicle (L1:50m, L2:75m, L3:100m).
6.2 Sim-to-Real Transfer
We evaluate the applicability of the CA-Opt attack in real driving environments by testing its effectiveness vs. baselines using a real-world dataset and on an actual vehicle.
Testing on a Real-world Dataset. We perform an evaluation using the comma2k19 dataset [61], a publicly available dataset with over 33 hours of California’s 280 highway commute. The dataset comprises 2019 segments, each lasting one minute, covering a 20km highway section, collected using OpenPilot hardware.
TABLE VIII: Performance of CA-Opt attack vs. CA-Random and CA-FGSM in deviating DNN-based lead vehicle position predictions using comma2k19 dataset.
Attack Metric Longitudinal Deviation in DNN Prediction (m)
0-20 20-40 40-60 60-80 80+ All
CA-Random Avg. 0.70 0.41 0.28 0.99 0.08 0.15
Std 0.69 1.03 1.44 1.93 2.67 2.56
CA-FGSM Avg. 6.78 4.08 4.47 2.15 0.90 1.29
Std 2.76 2.04 1.73 1.15 2.99 2.91
CA-Opt Avg. 18.65 16.15 14.52 8.65 3.73 4.91
Std 6.96 4.83 4.95 2.82 3.03 3.35
Table VIII compares the CA-Opt attack with two context-aware baselines, CA-Random and CA-FGSM, for different distances between the Ego and lead vehicles. CA-Random has an average deviation of 0.15m, which does not significantly impact ACC system outputs or cause hazards as the ACC system typically keeps a following distance larger than 4 meters in the absence of attacks. When the Ego vehicle is close to the lead vehicle (less than 20m), CA-FGSM shows a 6.78m average deviation, highlighting some level of vulnerability to gradient-based attacks. In contrast, CA-Opt achieves the highest deviation of 18.65m, showcasing the effectiveness of proposed objective function (Section 3.3.2) in generating optimal perturbations with substantial impact.
Test on Actual Vehicle. An evaluation of the proposed method on an actual vehicle is also conducted. We run two versions of the ACC control software (with or without the adversarial patch generation) on an Ego vehicle, parking behind a lead vehicle at different distances from 10m to 50m in a parking lot, and observe the changes in the DNN prediction of relative distance. To better observe the changes, we modified the code to present the distance prediction on the device monitor. Experimental results show that the proposed CA-Opt introduces an average deviation of 16.2m, with an example shown in Fig. 13. We refrain from conducting additional tests on inducing forward collisions in actual vehicles on the road, given the elevated risk to the safety of both the driver and the vehicle.
@@ 6.2 Sim-to-Real Transfer
Figure 13: Evaluation on an actual car: relative distance predictions with and without adding an adversarial patch.
6.3 Runtime Overhead
@@
To further evaluate the applicability of the proposed CA-Opt in real-world settings, we conduct tests to measure its runtime overhead on an actual device (Comma3) that runs ACC systems with the proposed adversarial patch generation. Specifically, we park the Ego vehicle in a parking lot behind a lead vehicle and enable ACC function with a cruise speed set to 0 mph. We record the time overhead for each component, as shown in Fig. 14, and report the average value over 5,000 control cycles (10ms per cycle).
@@
Experimental results show that the time overhead introduced by the context inference component before activating the attacks is minimal (1.17 us). Following the activation of attacks, the time overhead for the object detection module is about 10.1 ms on average. Note that some production ADAS provide object detection and tracking features, so this overhead time could be potentially avoided. We also observe that the primary attribution algorithm [48] does not add significant overhead, leveraging gradients calculated during the patch optimization process. The combined time overhead for primary attribution analysis, patch optimization, and the patch update algorithm is 1.52 ms in total.
Notably, the attacker refrains from modifying the initial image immediately after activating the attack to preserve the temporal relationship of the original process. Instead, the attacker awaits the outcomes of object detection and DNN perception applied to the first image before executing the adversarial patch generation algorithms. Subsequently, the generated patch is introduced to the first image, which is then inputted into the DNN model during the subsequent perception cycle. Consequently, this attack process introduces a delay equivalent to one perception cycle. Nevertheless, the advantage lies in its non-disruptive alteration of the original ACC execution sequence and the effective generation of the runtime perturbations.
@@
6.4 Evading Existing Defense Methods
While the proposed CA-Opt attack can create stealthy adversarial patches invisible to the human eye, it may be detected by existing defense methods.

### v3

@@ V-B Attack Success Rate in Causing Hazards
We consider an attack successful if a collision event is observed (Ego vehicle collides with the lead vehicle) or the relative distance between the lead vehicle and the Ego vehicle is no larger than 0m. The success rate is reported across 1000 simulations, if not specified otherwise.
Figure 9: Top: Adversarial patch examples generated using our optimization method vs. random and APGD-based methods. Bottom: Success rate of CA-Opt and baseline attacks in absence of safety interventions.
Fig. 9 displays the success rates for each attack. The CA-Random attack leads to hazards in less than 1.2% of scenarios, with an overall success rate of 0.7%. This low rate suggests that randomly generated adversarial patches minimally impact the DNN model predictions and rarely cause hazards. Increasing the random perturbation values in the adversarial patch also does not significantly enhance the success rate. This is due to the DNN model in OpenPilot being primarily trained to detect any front objects, not to identify their specific class, thus showing greater resilience to adversarial attacks. Additionally, patches with larger perturbation values are more visible to the driver compared to those from the optimization-based method (Fig. 9-Top), potentially alerting the driver to prevent hazards.
In contrast, the proposed CA-Opt attack achieves a 100% success rate in all testing scenarios, surpassing CA-Random by 142.9 times. Although CA-APGD uses a similar goal function as CA-Opt, it does not cause hazards in 46.6% of simulations, mainly due to limiting its number of iterations to satisfy real-time constraints. On the other hand, by employing a dynamic patch updating algorithm, our proposed CA-Opt attack ensures the continuity of the optimization process across the perception cycles and enhances the attack effectiveness.
We also observe that the Random-Opt baseline only achieves an average success rate of 3.5%, 28.6 times lower than the CA-Opt, as it wastes resources by injecting perturbations at non-critical system states. This highlights the importance of strategic timing of attacks and the insufficiency of optimization-based methods alone in causing hazards.
Observation 1: CA-Opt is more efficient than baselines in identifying the most critical times and optimal DNN perturbation values for attacking the ACC systems at runtime and overcoming real-time constraints (C3).
V-C Attack Stealthiness with Safety Interventions
@@
V-C1 Stealthiness in Perception Input
@@
Table VI presents the results averaged over all the test scenarios and simulations. The CA-Opt attack achieves at least a 99.2% success rate under all three stealthiness levels and keeps the perturbation degree less than 0.015 ($L_{\infty}$) and 0.184 ($L_{2}$). The perturbed image with the adversarial patch has a similarity of UIQ = 0.993 (1 means identical) to the original image. We choose the $\lambda$ value to be $10^{-3}$ in our evaluations because of its stealthiness and high attack effectiveness. Examples of the generated adversarial patches (with $\lambda=10^{-3}$) are presented in Fig. 5 (see the zoomed-in area) and Fig. 9, which are almost invisible to human eyes.
To further evaluate the stealthiness of our attack design, we also conduct a user study with 30 participants. Results show that adversarial patches at $\lambda=10^{-2}$ and $\lambda=10^{-3}$ are almost imperceptible to human drivers, and the patches generated by CA-Opt attacks are less noticeable than those generated by baseline perception attacks (CA-Random and CA-APGD) (refer to Appendix D for more details).
TABLE VI: Attack success rate with different patch stealthiness levels $\lambda$ and pixel perturbation degrees.
@@
V-C2 Evading Safety Interventions
For a more realistic evaluation of the effectiveness of different attack strategies, we rerun our experiments with different safety interventions (introduced in Section IV). A calibration of the safety features is performed before our evaluations to make sure the interventions are triggered correctly without any false positives.
We test each attack method with different AEBS configurations: (i) FCW/AEB depends on an independent camera that is not compromised, (ii) AEB/FCW utilizes compromised camera inputs similar to the ACC (simulating stock ACC and AEBS that share a camera or independent ACC and AEBS cameras that are both compromised), or (iii) AEB/FCW is disabled. Driver intervention and ACC safety constraint checking (OpenPilot Panda checks) are considered for all three settings. Here, we do not test the Random-Opt attack due to its similarity to CA-Opt but with worse performance. We assess the efficacy of each attack method using metrics such as the attack success rate, safety intervention activation rate (indicating the percentage of simulations triggering safety interventions), and hazard prevention rate (the percentage of simulations where hazards occur without safety interventions).
Table VII shows the experimental results of each attack method with different safety intervention configurations. We observe that, regardless of the interventions, the CA-Random and CA-APGD attacks fail to cause any hazards due to their low baseline success rates (see Fig. 9) and their noticeable perturbations that trigger the driver interventions in 23.8-27.4% and 100% of scenarios. These findings highlight the effectiveness of human drivers in preventing accidents and keeping autonomous driving safe.
In contrast, with AEBS disabled (the last row of Table VII), the CA-Opt attack successfully evades driver intervention in all simulations. This is attributed to the small, virtually invisible value of the adversarial patch, resulting in an average attack success rate of 100%. We also conduct experiments that simulate higher driver sensitivity levels by decreasing the mean perturbation value threshold for activating driver intervention from the default value of 15% (see Section 6) to 10%, 5%, 2%, 1.5%, 1%, and 0.5%. As shown in Fig. 10, when perturbation thresholds are set to 0.5% and 1% (representing highly sensitive driver), the CA-Opt attack triggers driver interventions in all and 79.6% of the simulations, leading to attack success rates of 0% and 20.4%, respectively. However, with thresholds higher than 1.5%, our attack maintains a 100% success rate without triggering driver intervention. This finding underscores the robustness of the attack in evading driver detection across a range of driver sensitivities.
TABLE VII: Performance of attacks with all the safety features and different AEBS settings.
@@ V-C2 Evading Safety Interventions
V-D Comparison to DNN Output and Control Output Attacks
The stealthy perturbations on the perception input can get propagated through the DNN model and ACC logic and lead to changes in the DNN output ( 2 in Fig. 1) and ACC control output 3. Although the attacker’s goal is to maximize errors in the DNN output and cause sudden accelerations on the ACC output, large deviations in vehicle states may be detected by the human driver or existing safety and defense mechanisms. To further evaluate the stealthiness of our proposed attack (CA-Opt), we compare the deviations resulting from the attack to those caused by stealthy attacks directly on DNN and control outputs. Note that such attacks are only possible under specific threat models (e.g., malware or wireless methods) in Table II.
Control Output Attacks. We first examine deviations in the autonomous vehicle states and control outputs resulting from the attack compared to two baseline output attacks, called MaxOut and StrategicOut. These attacks directly modify ACC output control commands, by setting them to a maximum allowed acceleration value (MaxOut) or a strategic value (StrategicOut) based on a method from prior research [97]. But they use the same context-aware method as CA-Opt for selecting the attack times and durations.
@@ V-D Comparison to DNN Output and Control Output Attacks
TABLE VIII: Performance of StrategicOut attack with all the safety features and different AEBS settings (AEBS with Shared Camera).
@@
We also evaluate the success rate of StrategicOut attack under two different safety intervention configurations. We do not assess the MaxOut attack due to its high likelihood of being detected. Table VIII shows that without AEBS, StrategicOut achieves a high success rate of 99.5% by generating attack values within safety limits and avoiding driver intervention. However, with AEBS active, using the same camera inputs as ACC, the success rate drops significantly to 20.3%, primarily due to safety interventions triggered in all simulations.
Figure 12: CA-Opt perception attack vs. DNN output attack.
Similarly, we compare the performance of CA-Opt attack with a stealthy attack that directly compromises DNN output 2 (referred to as DNNOut) by formulating an optimization problem to maximize the RD prediction within one standard deviation while ensuring the satisfaction of safety constraints on acceleration and speed [97]. We calculate the acceleration and speed values corresponding to RD predictions by replicating the Openpilot MPC and PID algorithms. This baseline uses the same context-aware method as CA-Opt for selecting the attack times and durations. We observe that the DNNOut attack causes a more obvious change in the RD predictions (see Fig. 12) compared to CA-Opt attack on DNN inputs 1, which then results in similar obvious changes in the gas, speed, or acceleration as depicted in Fig. 11.
Observation 4: CA-Opt attack has advantage over direct DNN or control output attacks in minimizing vehicle state changes to evade detection by safety interventions, while maintaining high effectiveness in causing hazards.
VI Evaluation in Real-World Settings
In this section we aim to answer the following questions about the effectiveness of our attack in real-world settings.
RQ4: Is our attack robust to real-world factors such as different camera positions and weather conditions?
RQ5: Can our attack transfer well from simulation to real-world implementation?
RQ6: How does our attack affect the real-time system operation?
RQ7: Can our attack evade detection or mitigation by the existing adversarial patch defense methods?
VI-A Robustness to Real-world Factors
@@
Fig. 13 illustrates the 100% success rate of our CA-Opt attack across 12 testing scenarios. The Ego vehicle initially maintains a safe following distance, deviates from it around the 2,500-3,000 control cycle or step due to the adversarial patch, and eventually collides with the lead vehicle. These results demonstrate our attack is robust to different camera positions and initial longitudinal distances and can cause safety hazards.
@@
Figure 13: Actual relative distance trajectories under CA-Opt attack with different camera heights (H1:1.1m, H2:1.3m, H3:1.5m, H4:1.7m) and initial longitudinal distances to the lead vehicle (L1:50m, L2:75m, L3:100m). An actual relative distance of zero indicates collision.
VI-B Performance on Actual Vehicles
We assess the feasibility of the CA-Opt attack on an actual vehicle (Lexus NX 2020) equipped with a production L2 ADAS, Comma 3, running OpenPilot software v0.8.9 by examining the attack impact on (i) the DNN perception module only and (ii) the end-to-end ACC.
Perception Module Evaluation. We evaluate the perception module in two scenarios of approaching an LV when (i) parked in a parking lot and (ii) driving on an actual road. In each scenario, the ACC on the Ego vehicle was tested with and without the adversarial patches injected to the camera frames.
First, the Ego vehicle was parked at distances ranging from 10m to 50m (at intervals of 5m) to the LV. We modified the OpenPilot code to display the RD predictions on the device monitor, as shown in Fig. 15-1a,1b. In these tests, the CA-Opt attack caused an average deviation of 16.2m in distance predictions, which could likely lead to a forward collision in the end-to-end ACC. This conclusion is based on our simulation experiments, where deviations exceeding 10 meters triggered sudden accelerations, leading to forward collisions.
@@ VI-B Performance on Actual Vehicles
Figure 14: Relative distance predictions with (solid lines) or w/o (dashed lines) adversarial patch for different driving scenarios.
Then we conducted experiments using the same scenarios (SC1-SC4) outlined in Section V-A. For an accurate assessment of the impact of the attack, we cloned the LVD’s DNN model within the OpenPilot control software and ran both the original and the duplicate model on the AV simultaneously. During each perception cycle (20Hz), we initially supplied a benign image (with an odd image index number) to the standard DNN model. Then, we duplicated this benign image, injected the adversarial patch to it, and then fed it to the second DNN model. The predictions from each model were recorded in separate log files. The results, presented in Figure 14, show that the attack significantly increased the relative distance predictions for 15.3m on average in all tested scenarios.
We also conducted experiments using a real-world video dataset as described in Appendix F. These experiments demonstrate the effectiveness of CA-Opt attack in impacting the DNN-based perception module in real-world driving scenarios.
Figure 15: RD predictions with (1a) or w/o (1b) patch on an actual vehicle in a parking lot; (2a) Side view of lead car model; (2b) AV under perception attack collides with the lead car model; (2c) AV follows the car model in a benign scenario; (2d) Driver’s view upon collision. [Anonymous video: https://sites.google.com/view/CAP-Attack]
End-to-End Evaluation. We also evaluate the impact of attacks on the end-to-end ACC on actual vehicle. To ensure the safety of both the driver and the vehicle, we constructed a lead car model from PVC pipe, designed to match the dimensions of a real BMW car model [16]. We aimed for the OpenPilot system to recognize this fabricated car as a genuine vehicle by attaching a rear-view image of a car to its rear end (see Fig. 15-2c). In this experiment, the AV approached the LV from a distance of 50 meters with a cruise speed set at 28 mph. Meanwhile, the LV was propelled by two remote-controlled ground robots (Fig. 15-2a). We conducted this experiment with and without the attack (adversarial patches) activated.
OpenPilot software successfully recognized the lead car model as a legitimate vehicle and maintained a safe following distance and speed (about 5 mph) in the benign scenario (see Fig. 15-2c). However, in the presence of the attack, we observed that the AV continued to advance toward the lead car model and eventually collided with it (Fig. 15-2b), despite the AEBS being activated (Fig. 15-2d). This underscores the generalization of our proposed attack in efficiently causing safety hazards and exposes the inadequacy of existing safety mechanisms in preventing the attack. Moreover, the time elapsed between the AEBS warning and the collision was approximately 1.5-2.2 seconds, shorter than the average driver reaction time of 2.5 seconds, leaving insufficient time for a human driver to intervene and prevent the collision.
VI-C Runtime Overhead
@@
To further assess our attack’s real-world applicability, we measured its runtime overhead on a Comma3 device. We parked the Ego vehicle, equipped with OpenPilot and our attack malware, behind a lead vehicle in a parking lot, activating the ACC function with a cruise speed set to 0 mph. We record the time overhead for each component, as shown in Fig. 16, and report the average value over 5,000 control cycles.
@@
Experimental results show that the time overhead introduced by the context inference component before activating the attacks is minimal (1.17 us). Following the activation of attacks, the time overhead for the object detection module is about 10.1 ms on average. Note that some production ADAS provide object detection and tracking features, so this overhead time could be potentially avoided. We also observe that the primary attribution algorithm [83] does not add significant overhead, leveraging gradients calculated during the patch optimization process. The total time overhead is 1.52 ms.
@@
VI-D Evading Existing Defense Methods
While the proposed CA-Opt attack can create stealthy adversarial patches invisible to the human eye, it may be detected by some existing defense methods.